OpenStudy (anonymous):

WILL FAN AND MEDAL!!!! What is the value of the x variable in the solution to the following system of equations? 4x + 8y = 8 x + 2y = 2 0 2 There is no x value as there is no solution to this system x can be any value as there are infinitely many solutions to this system

OpenStudy (anonymous):

x is 0 and y is 1

jimthompson5910 (jim_thompson5910):

solve the second equation for x x + 2y = 2 x = 2 - 2y

jimthompson5910 (jim_thompson5910):

then plug this into the first equation 4x + 8y = 8 4(2 - 2y) + 8y = 8 8 - 8y + 8y = 8 8 + 0y = 8 8 + 0 = 8 8 = 8 since that equation is always true, this means that there are infinitely many solutions

OpenStudy (anonymous):

Okay I just wanted to double check. Thanks!

jimthompson5910 (jim_thompson5910):

yw